OpenStudy (anonymous):

A typist for kelly services reports to 3E's properties for work at 10:00 Am and leaves at 6:00 pm after having typed from the end of page 8 to the end of the page 50 of a proposal. 3 E's pays 120$ for the typists services. A) find the rate of pay, in dollars per hour. b) find the average typing rate, in number of pages per hour. C) find the rate of pay, in dollars per page.

OpenStudy (anonymous):

A) 15 B) 5.25 C) 2.86

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

welcome:)

OpenStudy (ajprincess):

Number of hours between 10a.m and 6p.m=8 A)Rate of pay, in dollars per hour=Total pay/Number of hours =120/8 =15 B)Average typing rate=Number of pages typed/Number of hours =(50-8)/8 =42/8 =5.25 C)Rate of pay, in dollars per page=Total pay/Number of pages =120/(50-8) =120/42 =2.857 =2.86 Is this clear @ja7?

OpenStudy (anonymous):

thx yea

OpenStudy (ajprincess):

yw:)

OpenStudy (anonymous):

merci beaucoup.

OpenStudy (ajprincess):

u r welcome:)